Seven resistors, all the same value, are connected in series as shown . If the current in the circuit is 40 mA, what is the value of each resistor? (voltage= 14V)
1
Expert's answer
2020-05-21T13:45:52-0400

We know that each resistor carries the same current if the resistors are connected in series. Also, we know that if we connect n equal resistors in series, their total resistance R will be


"R=nr."

According to Ohm's law, find the resistance in terms of voltage and current:


"R=nr=\frac{V}{I},\rightarrow r=\frac{V}{nI},\\
r=\frac{14}{7\cdot0.04}=50\space\Omega."